OpenStudy (anonymous):

Add: (–4 + n – 2n^2) + (2 – 5n – 6n^2)

OpenStudy (anonymous):

-8 n^2-4 n-2

OpenStudy (anonymous):

Can you walk it through with me? I keep getting stuck.

OpenStudy (gudden):

@LegoMyEgo ... lemme help u.. first of all open the bracket

OpenStudy (anonymous):

–4 + n – 2n^2 + 2 – 5n – 6n^2

OpenStudy (gudden):

now bring the terms with n^2 together and those with n together and also keep the constant terms together

OpenStudy (anonymous):

8n^3?

OpenStudy (gudden):

let me do this step and u see where were u commiting the mistake

OpenStudy (gudden):

-6n^2 -2n^2 -5n +n -4 +2 what i have done is that arranged the terms with simillar power together

OpenStudy (anonymous):

-6n^2+2n^2-5n+n+2

OpenStudy (gudden):

u missed -4

OpenStudy (anonymous):

-6n^2+2n^2-5n+n+2 -4

OpenStudy (texaschic101):

-6n^2 - 2n^2 - 5n + n -4 + 2 combine like terms -6n^2 - 2n^2 = -8n^2 -5n + n = -4n -4 + 2 = -2 now put them all together and you get : -8n^2 -4n - 2

OpenStudy (gudden):

now solve the simillar terms i.e. terms which have same power..

OpenStudy (anonymous):

8n^2+5n^2+-2

OpenStudy (texaschic101):

where are you getting that from ?

OpenStudy (gudden):

check again...

OpenStudy (anonymous):

Where did I mess up?

OpenStudy (gudden):

u messed up while adding the terms with the power 1 i.e. terms with which have "n" at the end

OpenStudy (anonymous):

So I don't add them?

OpenStudy (gudden):

while adding those with n squre too

OpenStudy (texaschic101):

like terms are just like they say....alike. 2n + 3n are like terms. 2n^2 + 3n^2 are like terms 2n + 2n^2 are not like terms you can only add/subtract them of they are alike

OpenStudy (gudden):

what did u do..? i suggest u to take a look at texas's comment

OpenStudy (anonymous):

Oh I see

OpenStudy (anonymous):

I messed up on the like terms....so it would be like Texas said

OpenStudy (gudden):

yeah.. now lets solve it step by step.. first of all add the sums with power n^2

OpenStudy (gudden):

don't forget their signs...

OpenStudy (anonymous):

2n^2+2n^2

OpenStudy (anonymous):

(–4 + n – 2n^2) + (2 – 5n – 6n^2) = - 2n^2 - 6n^2 +n - 5n - 4 + 2 =-8n^2 - 4n - 2 (ANS.)

OpenStudy (gudden):

see.. the terms that have n^2 are -6n^2 and -2n?^2 so what would u get on adding them ??

OpenStudy (anonymous):

-8n^2

OpenStudy (anonymous):

simple adding guys (–4 + n – 2n^2) + (2 – 5n – 6n^2) = - 2n^2 - 6n^2 + n - 5n - 4 + 2 =- 8n^2 - 4n - 2 (ANS.)

OpenStudy (gudden):

yuhoo u are correct @LegoMyEgo ... now lets add those with n i.e. -5n and n add these

OpenStudy (anonymous):

-5n^2

OpenStudy (texaschic101):

way to go LegoMyEgo on the first part....but -5n + n = -4n

OpenStudy (gudden):

why would u write n^2 ??? we are adding not multipling.... and while adding u donot increase the power of the variable... try again...:D.. u would definitely do it..:D

OpenStudy (anonymous):

Oh yeah So now its -8n^2+-4n+2?

OpenStudy (anonymous):

-5n-n=-4n?

OpenStudy (texaschic101):

n is the same as 1n.....so -5n + 1n = -4n and then you have -4 + 2 which is -2.

OpenStudy (anonymous):

@LegoMyEgo see this (–4 + n – 2n^2) + (2 – 5n – 6n^2) = - 2n^2 - 6n^2 +n - 5n - 4 + 2 [opening the brackets] =-8n^2 - 4n - 2 (ANS.) [adding -2n^2 and -6n^2= -8n^2.] [adding n and -5n = -4n ] [adding -4 and 2= -2]

OpenStudy (anonymous):

Okay I see

OpenStudy (gudden):

no u are wrong while adding the constant term.. cos they are -4 + 2 and when u add them u are supposed to write that sign whose number was greater...

OpenStudy (anonymous):

I will just ask me teacher.

OpenStudy (texaschic101):

ask her what ? You do not understand what we just said ?

OpenStudy (gudden):

give it a try @LegoMyEgo ... don't loose hope...

OpenStudy (anonymous):

@LegoMyEgo i've solve u the whole ans. which thing u dont understand????

OpenStudy (anonymous):

I don't understand the last part

OpenStudy (gudden):

guys.. guys.. guys...!! don't be so rude...!! someday we even used to get confused with such questions... isn't it..? so be patient..

OpenStudy (anonymous):

Thank you @gudden

OpenStudy (gudden):

alright the last part then here u go..

OpenStudy (anonymous):

@LegoMyEgo when u add two things u put the sign of the bigger one to the ans when -4 and 2 are added as -4 is bigger if u dont consider the sign (4>2) so when we add both of them then we write the ans with -ive sign cz 4 has -ve sign. SO, -4 + 2 = -2

OpenStudy (gudden):

- 4 + 2 just solve it.. and tell me how do u solve it..

OpenStudy (anonymous):

-2

OpenStudy (gudden):

correct..:D

OpenStudy (anonymous):

Okay I get it now. I just got lost for a second. Thank you @gudden

OpenStudy (gudden):

you are welcome.. :D